OpenStudy (anonymous):

(system of equations) For a wedding, Shereda bought several rose bouquet and several carnation bouquets. The roses cost $15 per bouquet and the carnations cost $8 per bouquet. Shereda bought a total of 17 bouquets and paid $192. How many roses did she buy?

OpenStudy (tyteen4a03):

Set up two equations: (Let x be roses and y be carnations) x + y = 17 (Total number of bouquets) 15x + 8y = 192 (Total amount paid) You should be able to solve it from here. :)

OpenStudy (anonymous):

no actually im stuck were do i go off from there? do i use substitution or elimination?

OpenStudy (anonymous):

substitution could work, prob the easiest

OpenStudy (anonymous):

x=-y+17 15(-y+17)=192 ???

OpenStudy (anonymous):

15(-y+17)+8y=192 then solve for y then keep goin

OpenStudy (anonymous):

So y=9 ??

OpenStudy (anonymous):

so Shereda bought 9 carnation bouquets? and just do the same for x

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

im guessing you did that right i didnt do it lol

OpenStudy (anonymous):

but as long as you know how to do it you should get it right

OpenStudy (anonymous):

Thanks so much i'll double check as i go but i pretty much get it now

OpenStudy (anonymous):

alright :D

OpenStudy (anonymous):

oh and i ment x=9 and she bought 9 roses

OpenStudy (anonymous):

lol right! xD